OpenStudy (mayaal):

giving off a medal! Verify the identity: cot(x-pi/2)=-tan x

OpenStudy (anonymous):

hey rose :)

OpenStudy (anonymous):

okay bye :)

OpenStudy (mayaal):

Oh,I'm sorry I replied late.

OpenStudy (mayaal):

Can u plz help me @Iam_indian ?

OpenStudy (anonymous):

for u :)

OpenStudy (mayaal):

thnku:)

OpenStudy (mayaal):

do u know how to solve this prob?

OpenStudy (anonymous):

no :'(

OpenStudy (mayaal):

It's ok.Thanks anyway:)

OpenStudy (mayaal):

@zepdrix can u help me?

OpenStudy (mayaal):

@study100

OpenStudy (anonymous):

@neoh147 this awesome person got your back!

OpenStudy (mayaal):

hahaha=p

OpenStudy (anonymous):

cot(x-pi/2)=-tan x left hand side = cos(x-pi/2)/sin(x-pi/2) = [cos x cos (pi/2) + sin x sin (pi/2)]/ [sin x cos (pi/2) -sin (pi/2)cos x] = (0+sin x)/ (0 - cos x) =-(sin x/ cos x) = - tan x = right hand side (proven#) note that cos (pi/2)= 0

OpenStudy (mayaal):

thnku soooo much @neoh147 :)

OpenStudy (anonymous):

welcome @mayaal